# DreamVideo: High-Fidelity Image-to-Video Generation with Image Retention and Text Guidance

### 05 Dec 2023 v1

@@ Abstract
Image-to-video generation, which aims to generate a video starting from a given reference image, has drawn great attention. Existing methods try to extend pre-trained text-guided image diffusion models to image-guided video generation models. Nevertheless, these methods often result in either low fidelity or flickering over time due to their limitation to shallow image guidance and poor temporal consistency. To tackle these problems, we propose a high-fidelity image-to-video generation method by devising a frame retention branch on the basis of a pre-trained video diffusion model, named DreamVideo. Instead of integrating the reference image into the diffusion process in a semantic level, our DreamVideo perceives the reference image via convolution layers and concatenate the features with the noisy latents as model input. By this means, the details of the reference image can be preserved to the greatest extent. In addition, by incorporating double-condition classifier-free guidance, a single image can be directed to videos of different actions by providing varying prompt texts. This has significant implications for controllable video generation and holds broad application prospects. We conduct comprehensive experiments on the public dataset, both quantitative and qualitative results indicate that our method outperforms the state-of-the-art method. Especially for fidelity, our model has powerful image retention ability and result in high FVD in UCF101 compared to other image-to-video models. Also, precise control can be achieved by giving different text prompts. Further details and comprehensive results of our model will be presented in https://anonymous0769.github.io/DreamVideo/.
@@ 1 Introduction
Due to the large amount of potential applications in artistic creation, entertainment, etc., image or video generation has attracted much attention. Text-guided image generation has experienced explosive growth and a large number of excellent works have been published such as GLIDE [24], Imagen [9], Stable Diffusion [22], et al. Leveraging existing text-guided image synthesis technologies, text-to-video generation [26] has also achieved certain milestones. Text-to-video generation can be seen as a simple extension of text-to-image in video domain by injecting temporal layers into the network with spatial layers frozen. The ability of text-guided video generation is however limited in real applications. The reasons are twofold. One reason is that a simple and short sentence can barely describe a video containing dynamic content. The other reason is that complex texts present a significant challenge to the representational capabilities of the generation model, thereby greatly increasing the likelihood of generation failure. For this purpose, various methods have been proposed for generating videos guided by different conditions including optical flows [15], depth sequence [29], dragging strokes [33], et al. Among all the conditions, the condition of image is a natural and intuitive way for guiding video generation. On the one hand, an image provides substantial visual details. On the other hand, text prompts can still play a role for controlling the diversity of video generation. Therefore, image-to-video generation with text guidance holds significant importance.
@@
Image-to-video generation with text guidance is a straightforward idea but achieving high quality is nontrivial. The biggest challenge is how to enable flexible controlling over various video content with text guidance meanwhile maintaining the fidelity to the input image. Usually, visual details sacrifice for the temporal consistency across generated video frames, which leads to flickering over the generated video frames. For the same reason, all of the generated videos frames differ from the input one so the fidelity gets low. Moreover, even if both fidelity and temporal consistency can be maintained, the problem of limited capability of video motion control by text guidance, as another challenge, arises. Previous works have attempted to process the input image as an additional condition to text prompt [3]. The result is unsatisfactory especially in fidelity because visual information is only captured in semantic level. To tackle this problem, there are also works [35] using two stages training on images and then texts respectively by using cascaded diffusion models. The authors try to keep visual fidelity to the input image and meanwhile hold a good video-text alignment via two-stage training. Such complex design partially alleviate the problem of image-to-video generation we previously mentioned. However, the later refinement stage for text-guided video generation plays a greater role, even weakening the effects of the first image-guided stage. The result of this structure is that there is a challenging contradiction between fidelity and text alignment. This issue is also evident in the generation results of the I2VGen-XL [35].
@@
To tackle these problems, we propose a high-fidelity image-to-video generation method by devising a frame retention branch on the basis of a pre-trained video diffusion model. Instead of integrating the reference image into the diffusion process in a semantic level, our method perceives the reference image via convolution layers and concatenate the features with the noisy latents as model input. By this means, the details of the reference image can be preserved to the greatest extent. In addition, by incorporating double-condition classifier-free guidance, a single image can be directed to videos of different actions by providing varying prompt texts. Figure 1 exhibits three cases of image-to-video generation using DreamVideo, with the image examples sourced from MidJourney [17], which have high fidelity and great text alignment. Hence, our DreamVideo can seamlessly integrate with the text-to-image generation model, thereby facilitating a production pipeline from text to image to video. Moreover, for fully evaluating our proposed method, extensive experiments are conducted on multiple benchmarks. On the benchmark of image-video generation based on UCF101, we achieved a considerable Fréchet Video Distance (FVD) [28] score of 214.52, which significantly outperforms most existing methods. Also, the comparisons with other methods are conducted from multiple aspects including user study.
@@ 2.1 Video diffusion models
Diffusion Models (DMs) [8] show impressive results in image synthesis and lots of methods have been proposed such as GLIDE [16], Imagen [24], Stable Diffusion [22], et al. With the mature of image generation models, video generation also receives much attention and the widely adopted approach is injecting temporal layers into the image DMs to enable the network temporal representation. Lots of video diffusion models have been published such as Make-A-Video [26], CogVideo [10], Imagen Video [9], MagicVideo [36], VidRD [5], et al. For the training data, many works [1, 30] find that using both image and video data can significantly improve the appearance details and also alleviate the phenomenon of catastrophic forgetting. For the complexity of video data, PVDM [34] creatively proposes an image-like 2D latent space for efficient parameterization. Also, some works studies the effects of initial noise prior in video DMs. VideoFusion [14] finds the image priors of the pre-trained model can be efficiently shared by all frames and thereby facilitate the learning of video data. PYoCo [4] devises a video noise prior to get better temporal consistency. In addition, some works use additional DMs for frame interpolation, prediction and super resolution to achieve better performance. Align Your Latent [1] and LAVIE [30] are two relatively complete pipelines to generate high-quality videos. Both of them devise a basic video DM to generate the initial video frames and additional modules for termporal interpolation and Video Super Resolution (VSR).
@@ 2.2 Image-to-video generation
Earlier video generation works following image diffusion models are mostly text-guided. Though creative videos can indeed be generated with carefully devised text prompts, the appearance, layout or motion of the generated videos can not be precisely controlled. For this purpose, lots of works try to integrate other conditions or controls into video DMs. Image-to-video generation is a straightforward idea but not a trivial one to implement. AnimateDiff [6] injects motion modelling modules into frozen text-to-image models so motion priors can be distilled for different personalized models to implement image-to-video generation. VideoComposer [29] devises a unified interface for multiple conditions including single image condition. Also, there are some works exploiting additional modalities to generate videos from images. Dreampose [11] adopts sequence of poses and LFDM [15] uses optical flows. Additionally, DragNUWA [33] pushes image-to-video ahead by integrating user strokes to realize more controllable video generation. VideoCrafter1 [3] treats the input image as an additional condition and devises a dual attention module to encode and inject both text prompt and reference image. I2VGen-XL [35] uses two stages of cascaded diffusion models to achieve high semantic consistency and spatiotemporal continuity. However, most of these works on image-to-video generation are semantic level so the details of objects often differ from those in the input image. Therefore a high-fidelity image-to-video generation method is necessary.
@@ 4.1 Model Structure
Figure 2: The architecture of DreamVideo. A reference image is processed by a convolution block and concatenated with the representation of noisy latents. The Image Retention Module, as a side branch copying from the downsample blocks of U-Net, plays a role for maintaining the visual details from the input image and meanwhile also accepting text prompts for motion control.
The architecture of our model is shown in Fig. 2, which comprises of two principal module, the primary U-Net block and the Image Retention block. The primary U-Net block forms the foundation of VLDMs and is comprised of a Variational Auto-Encoder (VAE) [12] and U-Net. The VAE takes on the responsibility of encoding and decoding latent representations, whereas U-Net effectively handles denoising. The disparity between our U-Net and the one in Stable Diffusion [21] lies in the addition of Temp-Attn and Temp-Conv within each sample block in our model. These elements play a pivotal role in preserving the temporal coherence of the generated video. The Image Retention block encompasses the duplication of network parameters in every down-sample and middle layer of the primary U-Net. In addition, zero-convolutional layers are added at the end of each block to progressively increase the image retention control over the U-Net, which also can preserve the text-video generation ability of the primary U-Net. To facilitate the input of the video latent representation, the 2D convolutions within the network have been modified to 3D convolutions.
@@ 4.3 Classifier-free Guidance for Two Conditionings
Figure 3: These are the generated first frames using different image guidance scale (GS) for classifier-free guidance.
@@ 4.4 Inference
We propose a Two-Stage Inference method, which uses the final frame of a previously generated video as the initial frame for the subsequent video. This strategy enables us to extend the video length by feeding the same text into the Two-Stage Inference process. Moreover, we can generate varied content video by supplying different texts in two inference process. Thanks to the great image retention capability of our DreamVideo, we can make the generated video longer and get generated video with two plots by Two-Stage Inference method. The outcomes of our experiments using this method are presented in Sec. 5.5.
@@ 5.1 Experimental setup
Training datasets Since the quality of the generate video depends on the quality of the datasets, we select the high-quality public dataset Pexels 300K [18] for training data. This dataset contains videos with high resolution and strong diversity, particularly without any watermarks.
@@
    CLIP Similarity (CLIP-SIM). We compute the CLIP [20] text-image similarity for each individual frame and calculate the average score to the text prompt. Our evaluation utilizes the ViT-B/32 CLIP model as its backbone.
@@ 5.2 Qualitative evaluation
Figure 5: Given the same image, different text prompts can lead to different output videos. This can be seen as an evidence of our text-guidance capability for controling video content motion.
@@ 5.5 Two-Stage inference
Our model demonstrates a significant capacity for maintaining the initial frame in video outputs when give an initial control image, which allows for the exploration of a two-stage inference process. This process entails using the output video from the first stage as the initial image input for generating a video with preserved initial frame. Employing this two-stage inference enables our model to produce videos with up to 16 frames, maintaining continuity owing to its robust image retention capabilities. As depicted in the first row of Fig. 7, utilizing the same textual input across both stages results in an extended video length. Furthermore, as shown in the second row of Fig. 7, when different textual inputs are applied in each stage of inference, our model is capable of generating varied actions within a single video. Remarkably, despite these variations, the video retains consistency and a seamless flow.
@@ 6 Conclusion
In this work, we present DreamVideo, a model for synthesizing high-quality videos from images. Our DreamVideo has a great image retention capability and supports a combination of image and text inputs as controlling parameters. We propose an Image Retention block that combines control information and gradually integrates it into the primary U-Net. We Explore double-condition class-free guidance for different degrees of image retention. It’s noteworthy that one limitation of our model is that the image retention ability of our DreamVideo relies on high-quality training data. Our DreamVideo is enabled to generate video that maintains superior image retention quality through training with high quality datasets. Finally, we demonstrate DreamVideo’s superiority over open-source image-video model qualitatively and quantitatively.

### 10 Dec 2023 v2

@@ Abstract
Image-to-video generation, which aims to generate a video starting from a given reference image, has drawn great attention. Existing methods try to extend pre-trained text-guided image diffusion models to image-guided video generation models. Nevertheless, these methods often result in either low fidelity or flickering over time due to their limitation to shallow image guidance and poor temporal consistency. To tackle these problems, we propose a high-fidelity image-to-video generation method by devising a frame retention branch based on a pre-trained video diffusion model, named DreamVideo. Instead of integrating the reference image into the diffusion process at a semantic level, our DreamVideo perceives the reference image via convolution layers and concatenates the features with the noisy latents as model input. By this means, the details of the reference image can be preserved to the greatest extent. In addition, by incorporating double-condition classifier-free guidance, a single image can be directed to videos of different actions by providing varying prompt texts. This has significant implications for controllable video generation and holds broad application prospects. We conduct comprehensive experiments on the public dataset, and both quantitative and qualitative results indicate that our method outperforms the state-of-the-art method. Especially for fidelity, our model has powerful image retention ability and results in high FVD in UCF101 compared to other image-to-video models. Also, precise control can be achieved by giving different text prompts. Further details and comprehensive results of our model will be presented in https://anonymous0769.github.io/DreamVideo/.
@@ 1 Introduction
Due to the large amount of potential applications in artistic creation, entertainment, etc., image or video generation has attracted much attention. Text-guided image generation has experienced explosive growth and a large number of excellent works have been published such as GLIDE [24], Imagen [9], Stable Diffusion [22], et al. Leveraging existing text-guided image synthesis technologies, text-to-video generation [26] has also achieved certain milestones. Text-to-video generation can be seen as a simple extension of text-to-image in the video domain by injecting temporal layers into the network with spatial layers frozen. The ability of text-guided video generation is however limited in real applications. The reasons are twofold. One reason is that a simple and short sentence can barely describe a video containing dynamic content. The other reason is that complex texts present a significant challenge to the representational capabilities of the generation model, thereby greatly increasing the likelihood of generation failure. For this purpose, various methods have been proposed for generating videos guided by different conditions including optical flows [15], depth sequence [29], dragging strokes [33], et al. Among all the conditions, the condition of image is a natural and intuitive way to guide video generation. On the one hand, an image provides substantial visual details. On the other hand, text prompts can still play a role in controlling the diversity of video generation. Therefore, image-to-video generation with text guidance holds significant importance.
@@
Image-to-video generation with text guidance is a straightforward idea but achieving high quality is nontrivial. The biggest challenge is how to enable flexible control over various video content with text guidance while maintaining the fidelity to the input image. Usually, visual details are sacrificed for temporal consistency across generated video frames, which leads to flickering over the generated video frames. For the same reason, all of the generated video frames differ from the input one so the fidelity gets low. Moreover, even if both fidelity and temporal consistency can be maintained, the problem of the limited capability of video motion control by text guidance, as another challenge, arises. Previous works have attempted to process the input image as an additional condition to text prompt [3]. The result is unsatisfactory especially in fidelity because visual information is only captured at the semantic level. To tackle this problem, there are also works [35] using two-stage training on images and then texts respectively by using cascaded diffusion models. The authors try to keep visual fidelity to the input image and meanwhile hold a good video-text alignment via two-stage training. Such a complex design partially alleviates the problem of image-to-video generation we previously mentioned. However, the later refinement stage for text-guided video generation plays a greater role, even weakening the effects of the first image-guided stage. The result of this structure is that there is a challenging contradiction between fidelity and text alignment. This issue is also evident in the generation results of the I2VGen-XL [35].
@@
To tackle these problems, we propose a high-fidelity image-to-video generation method by devising a frame retention branch on the basis of a pre-trained video diffusion model. Instead of integrating the reference image into the diffusion process at a semantic level, our method perceives the reference image via convolution layers and concatenates the features with the noisy latents as model input. By this means, the details of the reference image can be preserved to the greatest extent. In addition, by incorporating double-condition classifier-free guidance, a single image can be directed to videos of different actions by providing varying prompt texts. Figure 1 exhibits three cases of image-to-video generation using DreamVideo, with the image examples sourced from MidJourney [17], which have high fidelity and great text alignment. Hence, our DreamVideo can seamlessly integrate with the text-to-image generation model, thereby facilitating a production pipeline from text to image to video. Moreover, to fully evaluate our proposed method, extensive experiments are conducted on multiple benchmarks. On the benchmark of image-video generation based on UCF101, we achieve a considerable Fréchet Video Distance (FVD) [28] score of 214.52, which significantly outperforms most existing methods. Also, the comparisons with other methods are conducted from multiple aspects including user study.
@@ 2.1 Video diffusion models
Diffusion Models (DMs) [8] show impressive results in image synthesis and lots of methods have been proposed such as GLIDE [16], Imagen [24], Stable Diffusion [22], et al. With the maturity of image generation models, video generation also receives much attention and the widely adopted approach is injecting temporal layers into the image DMs to enable the network temporal representation. Lots of video diffusion models have been published such as Make-A-Video [26], CogVideo [10], Imagen Video [9], MagicVideo [36], VidRD [5], et al. For the training data, many works [1, 30] find that using both image and video data can significantly improve the appearance details and also alleviate the phenomenon of catastrophic forgetting. For the complexity of video data, PVDM [34] creatively proposes an image-like 2D latent space for efficient parameterization. Also, some works study the effects of initial noise prior in video DMs. VideoFusion [14] finds the image priors of the pre-trained model can be efficiently shared by all frames and thereby facilitate the learning of video data. PYoCo [4] devises a video noise prior to getting better temporal consistency. In addition, some works use additional DMs for frame interpolation, prediction and super resolution to achieve better performance. Align Your Latent [1] and LAVIE [30] are two relatively complete pipelines to generate high-quality videos. Both of them devise a basic video DM to generate the initial video frames and additional modules for temporal interpolation and Video Super Resolution (VSR).
@@ 2.2 Image-to-video generation
Earlier video generation works following image diffusion models are mostly text-guided. Though creative videos can indeed be generated with carefully devised text prompts, the appearance, layout or motion of the generated videos can not be precisely controlled. For this purpose, lots of works try to integrate other conditions or controls into video DMs. Image-to-video generation is a straightforward idea but not a trivial one to implement. AnimateDiff [6] injects motion modeling modules into frozen text-to-image models so motion priors can be distilled for different personalized models to implement image-to-video generation. VideoComposer [29] devises a unified interface for multiple conditions including single image condition. Also, there are some works exploiting additional modalities to generate videos from images. Dreampose [11] adopts a sequence of poses and LFDM [15] uses optical flows. Additionally, DragNUWA [33] pushes image-to-video ahead by integrating user strokes to realize more controllable video generation. VideoCrafter1 [3] treats the input image as an additional condition and devises a dual attention module to encode and inject both text prompt and reference image. I2VGen-XL [35] uses two stages of cascaded diffusion models to achieve high semantic consistency and spatiotemporal continuity. However, most of these works on image-to-video generation are semantic level so the details of objects often differ from those in the input image. Therefore a high-fidelity image-to-video generation method is necessary.
@@ 4.1 Model Structure
Figure 2: The architecture of DreamVideo. A reference image is processed by a convolution block and concatenated with the representation of noisy latents. The Image Retention Module, as a side branch copying from the downsample blocks of U-Net, plays a role in maintaining the visual details from the input image and meanwhile also accepting text prompts for motion control.
The architecture of our model is shown in Fig. 2, which comprises two principal modules, the primary U-Net block and the Image Retention block. The primary U-Net block forms the foundation of VLDMs and is comprised of a Variational Auto-Encoder (VAE) [12] and U-Net. The VAE takes on the responsibility of encoding and decoding latent representations, whereas U-Net effectively handles denoising. The disparity between our U-Net and the one in Stable Diffusion [21] lies in the addition of Temp-Attn and Temp-Conv within each sample block in our model. These elements play a pivotal role in preserving the temporal coherence of the generated video. The Image Retention block encompasses the duplication of network parameters in every down-sample and middle layer of the primary U-Net. In addition, zero-convolutional layers are added at the end of each block to progressively increase the image retention control over the U-Net, which also can preserve the text-video generation ability of the primary U-Net. To facilitate the input of the video latent representation, the 2D convolutions within the network have been modified to 3D convolutions.
@@ 4.3 Classifier-free Guidance for Two Conditionings
Figure 3: These are the generated first frames using different image guidance scales (GS) for classifier-free guidance.
@@ 4.4 Inference
We propose a Two-Stage Inference method, which uses the final frame of a previously generated video as the initial frame for the subsequent video. This strategy enables us to extend the video length by feeding the same text into the Two-Stage Inference process. Moreover, we can generate varied content videos by supplying different texts in two inference processes. Thanks to the great image retention capability of our DreamVideo, we can make the generated video longer and get generated video with two plots by the Two-Stage Inference method. The outcomes of our experiments using this method are presented in Sec. 5.5.
@@ 5.1 Experimental setup
Training datasets Since the quality of the generated video depends on the quality of the datasets, we select the high-quality public dataset Pexels 300K [18] for training data. This dataset contains videos with high resolution and strong diversity, particularly without any watermarks.
@@
    CLIP Similarity (CLIP-SIM). We compute the CLIP [20] text-image similarity for each individual frame and calculate the average score for the text prompt. Our evaluation utilizes the ViT-B/32 CLIP model as its backbone.
@@ 5.2 Qualitative evaluation
Figure 5: Given the same image, different text prompts can lead to different output videos. This can be seen as evidence of our text-guidance capability for controlling video content motion.
@@ 5.5 Two-Stage inference
Our model demonstrates a significant capacity for maintaining the initial frame in video outputs when given an initial control image, which allows for the exploration of a two-stage inference process. This process entails using the output video from the first stage as the initial image input for generating a video with a preserved initial frame. Employing this two-stage inference enables our model to produce videos with up to 16 frames, maintaining continuity owing to its robust image retention capabilities. As depicted in the first row of Fig. 7, utilizing the same textual input across both stages results in an extended video length. Furthermore, as shown in the second row of Fig. 7, when different textual inputs are applied in each stage of inference, our model is capable of generating varied actions within a single video. Remarkably, despite these variations, the video retains consistency and a seamless flow.
@@ 6 Conclusion
In this work, we present DreamVideo, a model for synthesizing high-quality videos from images. Our DreamVideo has a great image retention capability and supports a combination of image and text inputs as controlling parameters. We propose an Image Retention block that combines control information and gradually integrates it into the primary U-Net. We Explore double-condition class-free guidance for different degrees of image retention. It’s noteworthy that one limitation of our model is that the image retention ability of our DreamVideo relies on high-quality training data. Our DreamVideo is enabled to generate video that maintains superior image retention quality through training with high quality datasets. Finally, we demonstrate DreamVideo’s superiority over the open-source image-video model qualitatively and quantitatively.